# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ 4.2 Differentiability and minimisers of covariance loss function
$F_{f}^{\prime}(x)(y)$ $=\sum_{i}f^{\prime}(x_{i})y_{i},$ $x\in\ell_{2}((-1,\infty)),\ y\in\ell^{2}(\mathbb{R}).$
Remark 4.10.
For a finite set of indices $i\in\{1,\ldots,l\}$, $l\in\mathbb{N}$, the convergence $(f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i})/y_{i}\rightarrow 0$ is uniform $i$. This implies that $\frac{1}{\lVert y\rVert}(\sum_{i}^{l}f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i})\rightarrow 0$, which implies differentiability of $F_{f}$ for finite-dimensional $\mathcal{H}$. In infinite dimensions, the convergence of each term is no longer uniform in $i$, as now $i\in\mathbb{N}$, and the previous sum need not converge to 0. Compared to a finite-dimensional setting, in the infinite-dimensional setting we therefore need to more assumptions of $f$ to assert its convergence. Hence we must restrict the class of spectral functions $\mathscr{F}$. In particular, we require additionally that $f$ has minimum 0 and a derivative which is Lipschitz at 0.
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. If $\mathcal{W}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$ is a subspace of finite dimension that contains $g(U+tV)$ for all $t\in\mathbb{R}$, then the restriction ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}:\mathcal{W}\rightarrow\mathbb{R}$ of $F_{f}\circ\Lambda$ to $\mathcal{W}$ satisfies ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)=F_{f}\circ\Lambda\circ g(U+tV)$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g$ is Gateaux differentiable in $U$ in direction $V$ if and only if ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g$ is. Hence, by the chain rule, e.g. [17, Section 1.4.1] or [5, Theorem 12.2.2], it suffices to show that ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable on all of $(\mathcal{W},\lVert\cdot\rVert_{L_{2}(\mathcal{H})})$ and that $g$ is Gateaux differentiable in $U$ in direction $V$ to be able to conclude the Gateaux differentiability of $J_{f}=F_{f}\circ\Lambda\circ g$ in $U$ in direction $V$.
We now use the finite-dimensional result of [19, Theorem 1.1] on differentiability of spectral functions to deduce Fréchet differentiability of ${\left.\kern-1.2pt(G\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ for certain $\mathcal{W}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$ and $G:\mathbb{R}^{\dim{\mathcal{H}}}\rightarrow\mathbb{R}$. This is done in Proposition 4.12 and Proposition 4.14. For this purpose, we introduce the following definition.
@@
Let $m\in\mathbb{N}\cup\{\infty\}$. A set $\Omega\subset\mathbb{R}^{m}$ is symmetric if $Px\in\Omega$ for every $x\in\Omega$ and every permutation $P:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$. If $\Omega\subset\mathbb{R}^{\mathbb{N}}$ is symmetric, then a function $G:\Omega\rightarrow\mathbb{R}$ is symmetric if $G(Px)=G(x)$ for every $x\in\Omega$ and every permutation $P:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$.
As an example of a symmetric set and symmetric function, consider respectively $\ell^{2}((-1,\infty])$ and $F_{f}$ from (16) for any $f\in\mathscr{F}$.
In analogy to the eigenvalue map $\Lambda:L_{2}(\mathcal{H})_{\mathbb{R}}\rightarrow\ell^{2}(\mathbb{R})$, we define $\Lambda^{m}:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\mathbb{R}^{m}$ for any $m$-dimensional subspace $\mathcal{Z}\subset\mathcal{H}$ to be any map that sends $X\in L_{2}(\mathcal{Z})_{\mathbb{R}}$ to its eigenvalue sequence, which we order in a nonincreasing way.
@@
Let $m\in\mathbb{N}$ and let the set $\Omega\subset\mathbb{R}^{m}$ be open and symmetric, and suppose that $G:\Omega\rightarrow\mathbb{R}$ is symmetric. Let $\mathcal{Z}\subset\mathcal{H}$ be finite-dimensional and let $X\in L_{2}(\mathcal{Z})_{\mathbb{R}}$ be such that $\Lambda^{m}(X)\in\Omega$. Then the function $G\circ\Lambda^{m}:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\mathbb{R}$ is Fréchet differentiable at $X$ if and only if $G$ is Fréchet differentiable at $\Lambda^{m}(X)\in\mathbb{R}^{m}$. In this case the Fréchet derivative of $G\circ\Lambda^{m}$ at $X$ is
$(G\circ\Lambda^{m})^{\prime}(X)=\sum_{i}G^{\prime}(\Lambda^{m}(X))_{i}e_{i}\otimes e_{i}\in L_{2}(\mathcal{Z}),$
@@
By definition of the Fréchet derivative, $(G\circ\Lambda^{m})^{\prime}(X)\in L_{2}(\mathcal{Z})_{\mathbb{R}}^{*}$. By the Riesz representation theorem, $L_{2}(\mathcal{Z})_{\mathbb{R}}^{*}\simeq L_{2}(\mathcal{Z})_{\mathbb{R}}$, and we consider $(G\circ\Lambda^{m})^{\prime}(X)$ as an element of $L_{2}(\mathcal{Z})_{\mathbb{R}}$.
As a consequence of Proposition 4.12, we can now deduce the Fréchet differentiability of ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ for symmetric functions $F$ and for a suitable subspace $\mathcal{W}$ of $L_{2}(\mathcal{H})_{\mathbb{R}}$.
@@
Let $\mathcal{Z}\subset\mathcal{H}$ be a finite-dimensional subspace. Let $\mathcal{W}\coloneqq\{X\in L_{2}(\mathcal{H})_{\mathbb{R}}:\ \operatorname{ran}{X}\subset\mathcal{Z}\}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$. Let $F:\ell^{2}(\mathbb{R})\rightarrow\mathbb{R}$ be a symmetric function and let $X\in\mathcal{W}$. Then $\ker{X}^{\perp}=\operatorname{ran}{X}$, and if $F$ is Fréchet differentiable in $\Lambda(X)$, then ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable in $X\in\mathcal{W}$. In this case, the Fréchet derivative is given by
@@
Let $F_{f}$ and $g$ be defined in (16) and (17). Then $J_{f}$ is Gateaux differentiable on $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, and for any $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, the Gateaux derivative at $U$ in direction $V$ is given by
@@
A coercive and Gateaux differentiable function has a global minimum, that can be found among its stationary points. The following lemma establishes the coercivity of $J_{f}$ over every finite-dimensional subspace of $\mathcal{H}$.
@@
Let $f\in\mathscr{F}$ and $\mathcal{V}\subset\mathcal{H}$ be finite-dimensional. Then $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, i.e. $J_{f}(U_{n})\rightarrow\infty$ whenever $U_{n}\rightarrow\infty$ in $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$. In particular, $J_{f}$ has a global minimum on $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, which can be found among the stationary points of $J_{f}$.
Unless $\mathcal{H}$ is finite-dimensional, the function $J_{f}$ is no longer coercive on all of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. To show this, we exploit the property that the finite-dimensional ranges of a sequence $U_{n}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ do not need to lie in the same finite-dimensional subspace of $\mathcal{H}$.
@@
Let $\mathcal{H}=\ell^{2}(\mathbb{R})$ and $r=1$. Furthermore, let $(e_{k})_{k}$ be the standard basis of $\mathcal{H}$ and put $\mathcal{C}_{\textup{pos}}e_{k}=k^{-\alpha}e_{k}$ for $\alpha>1$. Let $U_{n}t=n^{\beta}te_{n}$ for $t\in\mathbb{R}$, with $\beta>0$. Then $U_{n}U_{n}^{*}e_{k}=\delta_{n,k}n^{2\beta}e_{n}$. It follows that $\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{k}=\delta_{n,k}n^{-\alpha+2\beta}e_{n}$. Hence $\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}\rVert_{2}^{2}=\langle\mathcal{C}_{\textup{pos}}^{1/2}U_{n}U_{n}^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{n},e_{n}\rangle=n^{-\alpha+2\beta}$ which is bounded from above for $\alpha>2\beta$. Therefore $J_{f}(U_{n})$ is bounded in $n$, while $\lVert U_{n}\rVert=n^{\beta}\rightarrow\infty$.
In the proof of Theorem 4.18, coercivity on finite-dimensional subspaces of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the form $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, $\mathcal{V}\subset\mathcal{H}$ finite-dimensional, is enough to show the existence of a global minimizer of $J_{f}$, because all the stationary points lie in one such finite-dimensional subspace.
We can now state the main result of this section.

### v3

@@ 4.3 Differentiability and minimisers of covariance loss function
$F_{f}^{\prime}(x)(y)$ $=\sum_{i}f^{\prime}(x_{i})y_{i},$ $x\in\ell^{2}((-1,\infty)),\ y\in\ell^{2}(\mathbb{R}).$
Remark 4.12 (Necessity of assumptions on $\mathscr{F}$).
For a finite set of indices $i\in\{1,\ldots,l\}$, $l\in\mathbb{N}$, the convergence $(f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i})/y_{i}\rightarrow 0$ is uniform in $i$. This implies that $\frac{1}{\lVert y\rVert}(\sum_{i}^{l}f(x_{i}+y_{i})-f(x_{i})-f^{\prime}(x_{i})y_{i})\rightarrow 0$, which implies differentiability of $F_{f}$ for finite-dimensional $\mathcal{H}$. In infinite dimensions, the convergence of each term is no longer uniform in $i$, as now $i\in\mathbb{N}$, and the previous sum need not converge to 0. Compared to a finite-dimensional setting, in the infinite-dimensional setting we therefore need more assumptions on $f$ to obtain the desired convergence. Hence we restrict the function $f$ to the class of spectral functions $\mathscr{F}$ from (16a). In particular, we require additionally that $f$ has minimum 0 and a derivative which is Lipschitz at 0.
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. If $\mathcal{W}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$ is a subspace of finite dimension that contains $g(U+tV)$ for all $t\in\mathbb{R}$, then the restriction ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}:\mathcal{W}\rightarrow\mathbb{R}$ of $F_{f}\circ\Lambda$ to $\mathcal{W}$ satisfies ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)=F_{f}\circ\Lambda\circ g(U+tV)$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g$ is Gateaux differentiable at $U$ in the direction $V$ if and only if ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g$ is. Hence, by the chain rule, e.g. [16, Section 1.4.1] or [5, Theorem 12.2.2], it suffices to show that ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable on all of $(\mathcal{W},\lVert\cdot\rVert_{L_{2}(\mathcal{H})})$ and that $g$ is Gateaux differentiable at $U$ in the direction $V$ in order to show the Gateaux differentiability of $J_{f}=F_{f}\circ\Lambda\circ g$ at $U$ in the direction $V$.
We now use the finite-dimensional result of [20, Theorem 1.1] on differentiability of permutation invariant functions of spectra of symmetric matrices to deduce Fréchet differentiability of ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ for certain $\mathcal{W}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$ and $F:\mathbb{R}^{\dim{\mathcal{H}}}\rightarrow\mathbb{R}$. This is done in Proposition 4.16, using Proposition 4.14. For this purpose, we introduce the following definition.
@@
Let $m\in\mathbb{N}\cup\{\infty\}$. A set $\Omega\subset\mathbb{R}^{m}$ is symmetric if $Px\in\Omega$ for every $x\in\Omega$ and every permutation $P:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$. If $\Omega\subset\mathbb{R}^{m}$ is symmetric, then a function $\mathcal{G}:\Omega\rightarrow\mathbb{R}$ is symmetric if $\mathcal{G}(Px)=\mathcal{G}(x)$ for every $x\in\Omega$ and every permutation $P:\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$.
As an example of a symmetric set and symmetric function, consider respectively $\ell^{2}((-1,\infty))$ and $F_{f}$ from (18) for any $f\in\mathscr{F}$.
Recall from Section 1.5 the definition of the eigenvalue map $\Lambda^{m}:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\mathbb{R}^{m}$ for an $m$-dimensional subspace $\mathcal{Z}\subset\mathcal{H}$ and $m\in\mathbb{N}$. The ordering of eigenvalues given by $\Lambda^{m}$ is nonincreasing. The following result relates the Fréchet differentiability of $\mathcal{G}\circ\Lambda^{m}$ and $\mathcal{G}$ for symmetric functions $\mathcal{G}$.
@@
Let $m\in\mathbb{N}$ and let the set $\Omega\subset\mathbb{R}^{m}$ be open and symmetric, and suppose that $\mathcal{G}:\Omega\rightarrow\mathbb{R}$ is symmetric. Let $\mathcal{Z}\subset\mathcal{H}$ be $m$-dimensional and let $X\in L_{2}(\mathcal{Z})_{\mathbb{R}}$ be such that $\Lambda^{m}(X)\in\Omega$. Then the function $\mathcal{G}\circ\Lambda^{m}:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\mathbb{R}$ is Fréchet differentiable at $X$ if and only if $\mathcal{G}$ is Fréchet differentiable at $\Lambda^{m}(X)\in\mathbb{R}^{m}$. In this case the Fréchet derivative of $\mathcal{G}\circ\Lambda^{m}$ at $X$ is
$(\mathcal{G}\circ\Lambda^{m})^{\prime}(X)=\sum_{i}\mathcal{G}^{\prime}(\Lambda^{m}(X))_{i}e_{i}\otimes e_{i}\in L_{2}(\mathcal{Z}),$
@@
By definition of the Fréchet derivative, $(\mathcal{G}\circ\Lambda^{m})^{\prime}(X)\in L_{2}(\mathcal{Z})_{\mathbb{R}}^{*}$. By the Riesz representation theorem, $L_{2}(\mathcal{Z})_{\mathbb{R}}^{*}\simeq L_{2}(\mathcal{Z})_{\mathbb{R}}$, and we consider $(\mathcal{G}\circ\Lambda^{m})^{\prime}(X)$ as an element of $L_{2}(\mathcal{Z})_{\mathbb{R}}$.
As a consequence of Proposition 4.14, which is a result on Fréchet differentiability for symmetric functions of spectra of operators in $L_{2}(\mathcal{Z})$ for $\dim{\mathcal{Z}}<\infty$, we can now deduce the Fréchet differentiability of ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ for symmetric functions $F$ and suitable finite-dimensional subspaces $\mathcal{W}$ of $L_{2}(\mathcal{H})_{\mathbb{R}}$.
@@
Let $\mathcal{Z}\subset\mathcal{H}$ be a finite-dimensional subspace. Let $\mathcal{W}\coloneqq\{X\in L_{2}(\mathcal{H})_{\mathbb{R}}:\ \operatorname{ran}{X}\subset\mathcal{Z}\}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$. Let $F:\ell^{2}(\mathbb{R})\rightarrow\mathbb{R}$ be a symmetric function and let $X\in\mathcal{W}$. Then $\ker{X}^{\perp}=\operatorname{ran}{X}$, and if $F$ is Fréchet differentiable at $\Lambda(X)$, then ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}:\mathcal{W}\rightarrow\mathbb{R}$ is Fréchet differentiable at $X\in\mathcal{W}$. In this case, the Fréchet derivative is given by
@@
Remark 4.17.
The stated differentiability of ${\left.\kern-1.2pt(F\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}:\mathcal{W}\rightarrow\mathbb{R}$ holds with respect to the subspace topology on $\mathcal{W}$ inherited from $L_{2}(\mathcal{H})$. That is, we consider $\mathcal{W}$ as a Hilbert space with its $\lVert\cdot\rVert_{L_{2}(\mathcal{H})}$ norm.
@@
Let $F_{f}$, $g$, and $J_{f}$ be as defined in (18), (19), and (20) respectively. Then $J_{f}$ is Gateaux differentiable on $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, and for any $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, the Gateaux derivative at $U$ in the direction $V$ is given by
@@
A coercive and Gateaux differentiable function has a global minimum, that can be found among its stationary points. The following lemma establishes the coercivity of $J_{f}$ over every finite-dimensional subspace of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, that is, it establishes the coercivity of $J_{f}$ over $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ for every finite-dimensional subspace $\mathcal{V}$ of $\mathcal{H}$.
@@
Let $f\in\mathscr{F}$ and $\mathcal{V}\subset\mathcal{H}$ be finite-dimensional. Then $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, i.e. $J_{f}(U_{n})\rightarrow\infty$ whenever $\lVert U_{n}\rVert\rightarrow\infty$. In particular, $J_{f}$ has a global minimum on $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$, which can be found among the stationary points of $J_{f}$.
Unless $\mathcal{H}$ is finite-dimensional, the function $J_{f}$ is not coercive on all of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. To show this, we exploit the property that the finite-dimensional ranges of a sequence $U_{n}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ do not need to lie in the same finite-dimensional subspace of $\mathcal{H}$.
@@
Let $\mathcal{H}=\ell^{2}(\mathbb{R})$ and $r=1$. Furthermore, let $(e_{k})_{k}$ be the standard basis of $\mathcal{H}$ and put $\mathcal{C}_{\textup{pos}}e_{k}=k^{-\alpha}e_{k}$ for $\alpha>1$. Let $U_{m}t=m^{\beta}te_{m}$ for $t\in\mathbb{R}$, with $\beta>0$. Then $U_{m}U_{m}^{*}e_{k}=\delta_{m,k}m^{2\beta}e_{m}$. It follows that $\mathcal{C}_{\textup{pos}}^{1/2}U_{m}U_{m}^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{k}=\delta_{m,k}m^{-\alpha+2\beta}e_{m}$. Hence $\lVert\mathcal{C}_{\textup{pos}}^{1/2}U_{m}U_{m}^{*}\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}=\langle\mathcal{C}_{\textup{pos}}^{1/2}U_{m}U_{m}^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{m},e_{m}\rangle=m^{-\alpha+2\beta}$ which is bounded from above for $\alpha>2\beta$. Therefore, for $\alpha>2\beta$, $\lVert g(U_{m})\rVert$ is bounded in $m$ by the triangle inequality, while $\lVert U_{m}\rVert=m^{\beta}\rightarrow\infty$. We now argue that $J_{f}(U_{m})$ is bounded in $m$. Let $\gamma_{m}$ be the eigenvalue of largest magnitude among the eigenvalues of $g(U_{m})$. Lemma A.3 implies $\gamma_{m}=\lVert g(U_{m})\rVert$ is bounded in $m$ for $\alpha>2\beta$. Now, $f(\gamma)\leq f(\gamma_{m})+f(-\gamma_{m})$ for every eigenvalue $\gamma$ of $g(U_{m})$, because $xf^{\prime}(x)<0$ for $x\not=0$ implies that $f$ increases as $\lvert x\rvert$ increases. By Lemma 4.10, at most $n+r$ eigenvalues of $g(U_{m})$ are nonzero. Because $f(0)=0$ for $f\in\mathscr{F}$ in (16a), we conclude from (20), (18) and continuity of $f$ that $J_{f}(U_{m})\leq(n+r)(f(\gamma_{m})+f(-\gamma_{m}))$ is bounded in $m$.
In the proof of Theorem 4.21, coercivity on finite-dimensional subspaces of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ of the form $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ for finite-dimensional $\mathcal{V}\subset\mathcal{H}$ is enough to show the existence of a global minimizer of $J_{f}$, because all the stationary points lie in one such finite-dimensional subspace.
4.4 Optimal low-rank posterior covariance approximations
We can now state the solutions to Problem 4.3 and Problem 4.4.